40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 304 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  1806   Wed Jul 29 13:15:35 2009 ClaraUpdatePEMDents = Bad??

I was in the lab last night accelerometerizing and noticed some dents on the tubes that stick out horizontally from the MC2 optical chamber (sorry, I don't know what they're called or what they do). One of them is pretty big... I don't know if this is a problem, but it probably isn't a good thing. Photos below:

big_dent1.png

big_dent2.png

small_dent.png

This last one is a little hard to see... I was having trouble getting a good angle on it, but it's there. Not quite as significant as the first one though. (The first two pictures are of the same dent.)

  1805   Wed Jul 29 12:14:40 2009 peteUpdateComputersRCG work

Koji, Pete 

Yesterday, Jay brought over the IO box for megatron, and got it working.  We plan to firewall megatron this afternoon, with the help of Jay and Alex, so we can set up GDS there and play without worrying about breaking things.  In the meantime, we went to Wilson House to get some breakout boards so we can take transfer functions with the 785, for an ETMX controller.  We put in a sine wave, and all looks good on the auto-generated epics screens, with an "empty" system (no filters on). Next we'll load in filters and take transfer functions.

Unfortunately we promised to return the breakout boards by 1pm today.  This is because, according to denizens of Wilson House, Osamu "borrowed" all their breakout boards and these were the last two!  If we can't locate Osamu's cache, they expect to have more in a day or two.

Here is the transfer function of the through filter working at 16KHz sampling. It looks fine except for the fact that the dc gain is ~0.8. Koji is going to characterize the digital down sampling filter in order to try to compare with the generated code and the filter coefficients.


Attachment 1: TF090729_1.png
TF090729_1.png
Attachment 2: TF090729_1.png
TF090729_1.png
  1804   Wed Jul 29 12:00:49 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

For the past couple of days I have been trying to understand and perform Koji's method for impedance measurement using the Agilent 4395A Network Analyzer (without the impedance testing kit). I have made some headway, but I don't completely understand what's going on; here's what I've done so far.

I have made several transfer function measurements using the attached physical setup (ImpedanceTestingPhysicalSetup.png), after calibrating the setup by placing a 50 Ohm resistor in the place of the Z in the diagram. The responses of the various impedances I've measured are shown in the attached plot (ImpResponses.png). However, I'm having trouble figuring out how to convert these responses to impedances, so I tried to derive the relationship between the measured transfer function and the impedance by simplifying the diagram Koji drew on the board for me (attached, ImpedanceTestingSetup.png) to the attached circuit diagram (ImpedanceTestingCktDiagram.png), and finding the expected value of VA/VR. For the circuit diagram shown, the equation should be VA/VR = 2Z/(50+Z). 50 Ohms is good to use for calibration because the expected value of the transfer function for this impedance is 1 (0 dB).

So I used this relationship to find the expected response for the various impedances, and I also calculated the impedance from the actual measured responses. I've attached a plot of the measured (red) and expected (black) response (top) and impedance (bottom) for a 1 nF capacitor (1nF.png). The expected and measured plots don't really match up very well; if I add extra inductance (7.6 nH, plots shown in blue), the two plots match up a little better, but still don't match very well. I suspect that the difference may come from the fact that for my analysis, I treated the power splitter as if it were a simple node, and I think that's probably not very accurate.

Anyway, the point of all this is to eventually measure the impedance of the circuit I created on Friday, but I don't think I can really do that until I understand what is going on a little better.

Attachment 1: ImpedanceTestingPhysicalSetup.png
ImpedanceTestingPhysicalSetup.png
Attachment 2: ImpResponses.png
ImpResponses.png
Attachment 3: ImpedanceTestingSetup.png
ImpedanceTestingSetup.png
Attachment 4: ImpedanceTestingCktDiagram.png
ImpedanceTestingCktDiagram.png
Attachment 5: 1nF.png
1nF.png
  1803   Wed Jul 29 11:58:57 2009 AlbertoUpdatePSLMC not locking

This morning I found the Mode Cleaner unlocked.

I check the sliders for the mirrors bias and they have not changed. Also the OSEMs readbacks show no change in the optics positions.

I don't understand what's wrong because in the previous days, in this state of alignmanet, it could lock.

I tried to tweak a little bit the periscope to check whether it was a problem of beam matching but that didn't help the cavity to lock.

I don't want to change the periscpe alignment to much becasue I believe it is still good and I suspect that there is something else going on.

  1802   Wed Jul 29 11:15:06 2009 steveBureaucracySAFETYKevin receives safety training

Kevin Vigue, our high school summer student went through the 40m specific safety traning yesterday.

  1801   Tue Jul 28 18:32:21 2009 KojiUpdateCDSRCG work

Peter and Koji,

We are constructing a setup for the new 40m CDS using Realtime Code Generator (RCG).
We are trying to put simulated suspensions and test suspension controllers on a different processors of megatron
in order to create a virtual control feedback loop. Those CDS processes are communicating
each other via a shared memory, not via a reflective memory for now.

After some struggles with tremendous helps of Alex, we succeeded to have the communication between the two processes.
Also we succeeded to make the ADC/DAC cards recognized by megatoron, using the PCI express extension card replaced by Jay.
(This card runs multi PCI-X cards on the I/O chasis.)

Next steps:
- Establish a firewall between the 40m network and megatron (Remember this)
- Make DTT and other tools available at megatron
- Try virtual feedback control loops and characterize the performance
- Enable reflective memory functionalities on megatron
- Construct a hybrid system by the old/new CDSs
- Controllability tests using an interferometer


Note on MATLAB/SIMULINK
o Each cdsIPC should have a correct shared memory address spaced by 8 bytes. (i.e. 0x1000, 0x1008, 0x1010, ...)

Note on MEDM
o At the initial state, garbage (e.g. NaN) can be running all around the feedback loops. They are invisible as MEDM shows them as  "0.0000".
To escape from this state, we needed to disconnect all the feedback, say, by turning off the filters.

Note on I/O chasis
o We needed to pull all of the power plugs from megatron and the I/O chasis once so that we can activate
the PCI-e - PCI-X extension card. When it is succeeded, all (~30) LEDs turn to green.

  1800   Tue Jul 28 16:03:14 2009 ClaraUpdatePEMGuralp Seismometer cable pin diagram

I mapped out the corresponding pins on both ends of the Guralp seismometer cable. Here is the diagram:

guralp_pin_diagram.png

The circular 26-pin end of the cable (that plugs into the seismometer) is labeled as above. The other end (the 39-pin end) is not physically numbered, so I just came up with a numbering system. They are both pictured on the non-cable end of the connector. The colored circles indicate the pin pairs.

 

FROM JENNE, 30JULY2009:  the Dsub end is 37 pin, not 39.

  1799   Mon Jul 27 19:55:19 2009 KojiHowToIOOLens selection: plano-convex? or bi-convex?

Q. When should we use plano-convex lenses, and when should we use bi-convex?

As I had the same question from Jenne and Dmass in a month,
I just like to introduce a good summary about it.
Lens selection guide (Newport)
http://www.newport.com/Lens-Selection-Guide/140908/1033/catalog.aspx

At a first order, they have the same function.
Abberation (= non-ideal behavior of the lens) is the matter.

  1798   Mon Jul 27 17:48:44 2009 JenneUpdateoplevsETMY oplev is still down for the count

ETMY oplev is still out of order.  Hopefully I'll get it under control by tomorrow. 

  1797   Mon Jul 27 14:43:34 2009 ChrisUpdate Photodetectors

I found two ThorLabs PDA55 Si photodetectors that says detect visible light from DC to 10MHz that I'm going to use from now on.  I don't know how low of a frequency they will actually be good to.

  1796   Mon Jul 27 14:12:14 2009 ranaSummarySUSTM Coil Noise Spectra
Rob noticed that the ITMY DAC channels were saturating occassionally. Looking at the spectrum we can see why.
With an RMS of 10000 cts, the peak excursions sometimes cause saturations.

There's a lot of mechanical noise which is showing up on the ITM oplevs and then going to the mirror via
the oplev servo. We need to reduce the mechanical noise and/or modify the filters to compensate. The ITM
COIL_OUT RMS needs to be less than ~3000.
Attachment 1: Coils.pdf
Coils.pdf
  1795   Mon Jul 27 09:34:07 2009 steveSummaryIOOAligning the mode cleaner

Quote:

I set the MC back to its good alignment (June 21st) using this procedure. The trend of the OSEM values over the last 40 days and 40 nights is attached.

Then I aligned the periscope to that beam. This took some serious periscope knob action. Without WFS, the transmission went to 2.7 V and the reflection down to 0.6V.

Then I re-aligned the MC_REFL path as usual. The beam was far enough off that I had to also re-align onto the MC LSC PD as well as the MC REFL camera (~2 beam radii).

Beams are now close to their historical positions on Faraday and MC2. I then restored the PZT sliders to their April snapshot and the X-arm locked.

Steve - please recenter the iris which is on the periscope. It has been way off for a long time.

So it looks OK now. The main point here is that we can trust the MC OSEMs.

Afterwards I rebooted c1susvme1 and c1susvme2 because they were skewed.

 

 I'm impressed by Rana's simple way to align the MC. IFO arms are locked or flashing. 20 days trend attached.

 

Attachment 1: 20dtrend.jpg
20dtrend.jpg
  1794   Sun Jul 26 16:05:17 2009 AlbertoUpdatePSLAligning the mode cleaner

Quote:

I set the MC back to its good alignment (June 21st) using this procedure. The trend of the OSEM values over the last 40 days and 40 nights is attached.

Then I aligned the periscope to that beam. This took some serious periscope knob action. Without WFS, the transmission went to 2.7 V and the reflection down to 0.6V.

Then I re-aligned the MC_REFL path as usual. The beam was far enough off that I had to also re-align onto the MC LSC PD as well as the MC REFL camera (~2 beam radii).

Beams are now close to their historical positions on Faraday and MC2. I then restored the PZT sliders to their April snapshot and the X-arm locked.

Steve - please recenter the iris which is on the periscope. It has been way off for a long time.

So it looks OK now. The main point here is that we can trust the MC OSEMs.

Afterwards I rebooted c1susvme1 and c1susvme2 because they were skewed.

 

 It is really surprising that we now have again the data from the MC OSEMs since up to two days ago the record looked corrupted (see the attachments in my entry 1774).

The reason I ended up severely misaligning the the MC is exactly that there wasn't anymore a reference position that I could go back to and I had to use the camera looking a the Faraday.

  1793   Sun Jul 26 13:19:54 2009 ranaUpdatePSLAligning the mode cleaner

I set the MC back to its good alignment (June 21st) using this procedure. The trend of the OSEM values over the last 40 days and 40 nights is attached.

Then I aligned the periscope to that beam. This took some serious periscope knob action. Without WFS, the transmission went to 2.7 V and the reflection down to 0.6V.

Then I re-aligned the MC_REFL path as usual. The beam was far enough off that I had to also re-align onto the MC LSC PD as well as the MC REFL camera (~2 beam radii).

Beams are now close to their historical positions on Faraday and MC2. I then restored the PZT sliders to their April snapshot and the X-arm locked.

Steve - please recenter the iris which is on the periscope. It has been way off for a long time.

So it looks OK now. The main point here is that we can trust the MC OSEMs.

Afterwards I rebooted c1susvme1 and c1susvme2 because they were skewed.

 

Attachment 1: Untitled.png
Untitled.png
  1792   Sat Jul 25 19:04:01 2009 KojiUpdatePSLAligning the beam to the Faraday

Quote:

Quote:

When I turned them on, the control signal in Pitch from WFS2 started going up with no stop. It was like the integrator in the loop was fed with a DC bias. The effect of that was to misalign the MC cavity from the good state in which it was with the only length control on (that is, transmission ~2.7, reflection ~ 0.4).

I don't know why that is happening. To exclude that it was due to a computer problem I first burtrestored C1IOO to July the 18th, but since that did not help, I even restarted it. Also that didn't solve the problem.

 

 

At least one problem is the mis-centering of the resonant spot on MC2, which can be viewed with the video monitors.  It's very far from the center of the optic, which causes length-to-angle coupling that makes the mulitple servos which actuate on MC2 (MCL, WFS, local damping) fight each other and go unstable.

I played with the MC alignment for the beam centering. After that, I restored the alignment values.



In principle, one can select the MC2 spot as one likes, while the transmitted beam axis to the IFO is not changed
as far as you are at the best alignment. This principle is almost trivial because the beam axis matches
to the input beam axis at the best alignment.
The alignment solution is not unique for a triangle cavity if we don't fix the end spot position.

In practice, this cruising of the MC2 spot is accomplished by the following procedure:
0) Assume that you are initially at the best alignment (=max transmission).
1) Slightly tilt the MC2.
2) Adjust MC1/MC3 so that the best transmission is restored.

I started from the following initial state of the alignment sliders:

BEFORE TRIAL

MC1 Pitch  +3.6242
MC1 Yaw  -0.8640
MC2 Pitch  3.6565
MC2 Yaw -1.1216
MC3 Pitch -0.6188
MC3 Yaw -3.1910
MC Trans 2.70

After many iterations, the spot was centered in some extent. (See the picture)
RESULT

    adj.
MC1 Pitch  +3.363 (-0.26)
MC1 Yaw  -1.164 (-0.3)
MC2 Pitch  3.7565 (+0.1)
MC2 Yaw -1.2800 (~ -0.16)
MC3 Pitch -0.841 (~ -0.22)
MC3 Yaw -3.482 (~ -0.29)
MC Trans 2.75  

The instability looked cured somewhat.
Further adjustment caused a high freq (10Hz at the camera) instability and the IMCR shift issue.
So I returned to the last stable setting.

Side effect:
Of course, if you move MC1, the reflected spot got shifted.
The spot has been apparently off-centered from the IMCR camera. (up and right)
At this stage, I could not determine what is the good state.
So, I restored the alignment of the MC as it was.
But now Alberto can see which mirror do we have to move in which direction and how much.

Attachment 1: MC2_Cam.jpg
MC2_Cam.jpg
  1791   Sat Jul 25 16:04:32 2009 robUpdatePSLAligning the beam to the Faraday

Quote:

When I turned them on, the control signal in Pitch from WFS2 started going up with no stop. It was like the integrator in the loop was fed with a DC bias. The effect of that was to misalign the MC cavity from the good state in which it was with the only length control on (that is, transmission ~2.7, reflection ~ 0.4).

I don't know why that is happening. To exclude that it was due to a computer problem I first burtrestored C1IOO to July the 18th, but since that did not help, I even restarted it. Also that didn't solve the problem.

 

 

At least one problem is the mis-centering of the resonant spot on MC2, which can be viewed with the video monitors.  It's very far from the center of the optic, which causes length-to-angle coupling that makes the mulitple servos which actuate on MC2 (MCL, WFS, local damping) fight each other and go unstable.

  1790   Sat Jul 25 13:49:28 2009 KojiUpdateGeneralMultiply Resonant EOM Update

Quote:

After speaking with Rana and realizing that it would be better to use smaller inductances in the flying-component circuit (and after a lot of tinkering with the original), I rebuilt the circuit, removing all of the resistors (to simplify it) and making the necessary inductance and capacitance changes. A picture of the circuit is attached, as is a circuit diagram.

A plot of the measured and simulated transfer functions is also attached; the general shape matches between the two, and the resonant frequencies are roughly correct (they should be 11, 29.5, and 55 MHz). The gain at the 55 MHz peak is lower than the other two peaks (I'd like them all to be roughly the same). I currently have no idea what the impedance is doing, but I'm certain it is not 50 Ohms at the resonant peaks, because there are no resistors in the circuit to correct the impedance. Next, I'll have to add the resistors and see what happens.

Stephanie, 

This is a quite nice measurement. Much better than the previous one.

1) For further steps, I think now you need to connect the real EOM at the end in order to incorporate
the capacitance and the loss (=resistance) of the EOM. Then you have to measure the input impedance
of the circuit. You can measure the impedance of the device at Wilson house.
(I can come with you in order to consult with Rich, if you like)

Before that you may be able to do a preparatory measurement which can be less precise than the Wilson one,
but still useful. You can measure the transfer function of the voltage across the input of this circuit,
and can convert it to the impedance. The calibration will be needed by connecting a 50Ohm resister
on the network analyzer.

2) I wonder why the model transfer function (TF) has slow phase changes at the resonance.
Is there any implicit resistances took into account in the model?

If the circuit model is formed only by reactive devices (Cs and Ls), the whole circuit has no place to dissipate (= no loss).
This means that the impedance goes infinity and zero, at the resonance and the anti-resonance, respectively.
This leads a sharp flip of the phase at these resonances and anti-resonances.

The real circuit has small losses everywhere. So, the slow phase change is reasonable.

  1789   Sat Jul 25 13:34:58 2009 KojiUpdateGeneralWeek 5/6 Update

Quote:

The last week I've started setting up the HeNe laser on the PSL table and doing some basic measurements (Beam waist, etc) with the beam scan, shown on the graph.  Today I moved a few steering mirrors that steve showed me from at table on the NW corner to the PSL table.  The goal setup is shown below, based on the UCSD setup.  Also, I found something that confused me in the EUCLID setup, a  pair of quarter wave plates in the arm of their interferometer, so I've been working out how they organized that to get the results that they did.  I also finished calculating the shot noise levels in the basic and UCSD models, and those are also shown below (at 633nm, 4mw) where the two phase-shifted elements (green/red) are the UCSD outputs, in quadrature (the legend is difficult to read).

 

 

Chris,

Some comments:

0. Probably, you are working on the SP table, not on the PSL table.

1. The profile measurement looks very nice.

2. You can simplify the optical layout if you consider the following issues
  A. The matching lenses just after the laser:
      You can make a collimated beam only with a single lens, instead of two.
      Just put a lens of f0 with distance of f0 from the waist. (Just like Geometrical Optics to make a parallel-going beam.)

      Or even you don't need any lens. In this case, whole optical setup should be smaller so that your beam
      can be accomodated by the aperture of your optics. But that's adequately possible.

  B. The steering mirrors after the laser:
      If you have a well elevated beam from the table (3~4 inches), you can omit two steering mirrors.
      If you have a laser beam whose tilte can not be corrected by the laser mount, you can add a mirror to fix it.

  C. The steering mirrors in the arms:
      You don't need the steering mirrors in the arms as all d.o.f. of the Michelson alignment can be adjusted
      by the beamsplitter and the mirror at the reflected arm. Also The arm can be much shorter (5~6 inches?)

  D. The lenses and the mirrors after the PBS:
      You can put one of the lenses before the PBS, instead of two after the lens.
      You can omit the mirror at the reflection side of the PBS as the PBS mount should have alignment adjustment.

The simpler, the faster and the easier to work with!
Cheers.

  1788   Fri Jul 24 21:02:46 2009 AlbertoUpdatePSLAligning the beam to the Faraday

This afternoon I kept working on the alignment of the beam so that it matches at the same the PSL periscope, the Mode Cleaner and the Faraday isolator at the input of the IFO.

The camera looking at the Farady showed a beam quite low from the center of the Faraday's entrance. I wanted to move it up.

After working on the periscope alignment and on the MC mirrors, I think I managed to moved it up a bit. To know whether that was enough or not I wanted to evaluate the alignment to the X arm by checking the value of TRX.

In order for the MC to be finely matched to the input beam from the periscope, the WFS controls have to be on. Before turning them on, I centered the beam on their QPDs and run the WFS_zero_offset script.

When I turned them on, the control signal in Pitch from WFS2 started going up with no stop. It was like the integrator in the loop was fed with a DC bias. The effect of that was to misalign the MC cavity from the good state in which it was with the only length control on (that is, transmission ~2.7, reflection ~ 0.4).

I don't know why that is happening. To exclude that it was due to a computer problem I first burtrestored C1IOO to July the 18th, but since that did not help, I even restarted it. Also that didn't solve the problem.

 

Flashes at ETMX show at least that the beam is going through the Farady. How well, I can't tell untill the MC is under full control.

 

I have to leave the lab now, but I can be back tomorrow to keep working on that.

  1787   Fri Jul 24 17:47:52 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

After speaking with Rana and realizing that it would be better to use smaller inductances in the flying-component circuit (and after a lot of tinkering with the original), I rebuilt the circuit, removing all of the resistors (to simplify it) and making the necessary inductance and capacitance changes. A picture of the circuit is attached, as is a circuit diagram.

A plot of the measured and simulated transfer functions is also attached; the general shape matches between the two, and the resonant frequencies are roughly correct (they should be 11, 29.5, and 55 MHz). The gain at the 55 MHz peak is lower than the other two peaks (I'd like them all to be roughly the same). I currently have no idea what the impedance is doing, but I'm certain it is not 50 Ohms at the resonant peaks, because there are no resistors in the circuit to correct the impedance. Next, I'll have to add the resistors and see what happens.

Attachment 1: BuiltCkt2_Picture_Simplified.png
BuiltCkt2_Picture_Simplified.png
Attachment 2: BuiltCkt2_Simplified.png
BuiltCkt2_Simplified.png
Attachment 3: Simplified.png
Simplified.png
  1786   Fri Jul 24 17:20:48 2009 JenneUpdateoplevsETMY oplev is iffy

ETMY oplev is currently a work in progress.  The HeNe beam is hitting the photodiode, but the spot size there is pretty much the size of the entire QPD.  Thus, the ETMY oplev isn't really useful right now.  I'm re-figuring things out (note to self: close to the laser, you have to use Gaussian optics...regular ray tracing doesn't really work), and hopefully will have the oplev back under control by the time Alberto is finished realigning the IFO, so this doesn't keep anyone from doing any exciting locking work.

  1785   Fri Jul 24 11:04:11 2009 AlbertoConfigurationComputerselog restarted

This morning I found the elog down. I restarted it using the procedure in the wiki.

  1784   Thu Jul 23 20:30:23 2009 AlbertoUpdatePSLBeam aligned to the Farady

After yesterday's changes in the MC cavity state today it was necessary to optimize the alignment to the Faraday.

The way I did it was by tuning the PSL periscope in pitch and yaw trying to maximize TRX with the arm locked. After a small change in either one of the two directions I first maximized the MC transmitted power and then I ran the alignment script for the X arm.

I explored the space for both pitch and yaw and the max that I could get from TRX was 0.91. I'm not sure whether the increase in TRX is entirely due to a better alignment to the Farady rather than to a higher MC transmitted power.

Also I'm not sure I'm well interpreting the image from the camera pointing at the Farady. I guess I need someone more familiar with it to tell me if it shows any sign of clipping.

Anyway, last week, even before the MC got misaligned, TRX didn't go above 0.90. So now I wonder whether it's the MC's fault or something else's if we have that value..

  1783   Thu Jul 23 10:05:38 2009 AlbertoUpdatePSLSummary of the latest adventures with the alignment of the mode cleaner

Alberto, Koji,

Summary of the latest adventures with the alignment of the mode cleaner

Prior events.

  • Last week, on July 12th the RFM network crashed (elog entry 1736). I don't know for sure which one was the cause and which one the effect, but also C1DAQADW was down and it didn't want to restart. Alex fixed it the day after.
  • On the evening of Sunday July 20th I noticed that the mode cleaner was unlocked. A closer inspection showed me that MCL was frozen at -32768 counts. To fix that I rebooted C1DCUEPICS and burtrestored to snapshots from the day before.
  • On Tuesday July 21st another failure of the RFM Network made necessary a reboot of the frame builder and of all front end computers (entry 1772). As a consequence, the mode cleaner couldn't get locked anymore, even if the mirror's sliders in the MC-Align MEDM screen were in the proper positions. At that time I missed to check the MC suspension positions as a way to ensure that the MC hadn't really changed. Although later, as it turned out, that would have been useless anyway since all the data record prior to the computers crash of that day somehow had been corrupted (entry 1774). Neither the MC2 LSC control or MC ASC control could engage so I (erroneously) thought that some tune of the periscope might help. So I did but, since the Mode Cleaner was misaligned, that had the effect of spoiling the good matching of the periscope to the MC cavity.
  • Yesterday, Wednesday July 22nd, I found out about the sticky slider effect (entry 1776). At that point we didn't have anymore a way to know that the MC optics were actually in their proper original alignment state because of the lack of a reference for those in the data record (as I wrote above). I had to go back to the periscope and fix the alignment.


Chronicles of periscope and MC alignment

Yesterday morning I started aligning the periscope but it turned out to be trickier than usual. With the ASC (Alignment Sensing Control) off and only the length controls on, the Mode Cleaner didn't lock easily, although I knew I wasn't very far from the sweet spot.

In the afternoon the struggle continued and the matching of the the beam to the MC cavity became just worse. At some point I noticed that the ASC inputs somehow had got on - although the ASC still looked disabled from the MClock MEDM main screen. So I was actually working against the Wave Front Sensors and further worsening the periscope alignment.

That hurled me to the weeds. After hours of rowing across the stormy waters of a four-dimensional universe I got to have occasional TEM00 flashes at the transmission but still, surprisingly, no MC locking. Confused, I kept tuning the periscope but that just kicked me off road again.

Then at about 7pm Koji came to my rescue and suggested a more clever and systematic way to solve the problems. He suggested to keep record of the MC mirrors alignment state and re-align the cavity to the periscope. Then we would gradually bring the cavity back to the original good position changing the periscope alignment
at the same time.

 

That would have worked straight away, if we hadn't been fighting against a subtle and cruel enemy: the 40m computer network. But I (as John Connor), and Koji (as the Terminator) didn't pull back.

Here's a short list of the kinds of weapons that the computers threw to us:

  1. After a while the FSS entered a funny state. It showed transmission: we had light at the MC (and even flashes) but the MEDM readout of the FSS transmitted power after the cavity was low (~0.019). Also the spot on the monitor showed a slightly different pattern from how I remembered it. On the other side the transmission camera didn't show that typical halo as usual.
  2. MCL was frozen at 32768. I ran the MCDown and MCUp script a couple of times and that unstuck it.
  3. On op340m we found that the MC autolocker script wasn't running. So I restarted it. Still nothing changed: bright and sharp flashes appeared on the monitor (sign of a not too bad alignment) but no lock.
  4. I rebooted C1IOO. No change.
  5. I rebooted C1DCUEPICS and burtrestored the EPICS computers to Jul 19th. No change.
  6. Then I burtrestored the c1psl.snapshot and that finally did something. The FSS reflected spot changed and the halo appeared again at its transmission camera. Soon after the MC got locked.


We then proceeded with Koji's plan. In an iterative process, we aligned the MC cavity maximizing the transmission and tuned the periscope in order to match the Faraday input of the interferometer. The last thing we did it by looking at the camera pointing at the Faraday isolator.

We found that we didn't have to tune the periscope much. That means that all afternoon I didn't really go too far, but the autolocker wasn't working properly, or it wasn't working at all.

Then we ran the alignment script for the X arm but it didn't work before we aligned the steering mirrors.

Then we ran it three times but could not get more than 0.87 at TRX. That means that there we still have to work on the alignment to the Faraday. That's job for today in the trenches of the lab.

 

  1782   Thu Jul 23 07:34:45 2009 AidanUpdateCDSAdded C2 MEDM screens to 40m SVN.

 

See Adhikari eLOG entry: http://nodus.ligo.caltech.edu:8080/AdhikariLab/194

  1781   Wed Jul 22 20:11:26 2009 peteUpdateComputersRCG front end

I compiled and ran a simple (i.e. empty) front end controller on scipe12 at wilson house.  I hooked a signal into the ADC and watched it in the auto-generated medm screens. 

There were a couple of gotchas:

1. Add an entry SYS to the file /etc/rc.local, to the /etc/setup_shmem.rtl line, where the system file is SYS.mdl.

2. If necessary, do a BURT restore.  Or in the case of a mockup set the BURT Restore bit (in SYS_GDS_TP.adl) to 1.

 

  1780   Wed Jul 22 18:04:14 2009 robOmnistructureComputersweird noise coming from Gigabit switch

in the rack next to the printer.  It sounds like a fan is hitting something.

  1779   Wed Jul 22 16:15:52 2009 Chris ZimmermanUpdateGeneralWeek 5/6 Update

The last week I've started setting up the HeNe laser on the PSL table and doing some basic measurements (Beam waist, etc) with the beam scan, shown on the graph.  Today I moved a few steering mirrors that steve showed me from at table on the NW corner to the PSL table.  The goal setup is shown below, based on the UCSD setup.  Also, I found something that confused me in the EUCLID setup, a  pair of quarter wave plates in the arm of their interferometer, so I've been working out how they organized that to get the results that they did.  I also finished calculating the shot noise levels in the basic and UCSD models, and those are also shown below (at 633nm, 4mw) where the two phase-shifted elements (green/red) are the UCSD outputs, in quadrature (the legend is difficult to read).

 

 

Attachment 1: Beam_Scan.jpg
Beam_Scan.jpg
Attachment 2: Long_Range_Michelson_Setup_1_-_Actual.png
Long_Range_Michelson_Setup_1_-_Actual.png
Attachment 3: NSD_Displacement.png
NSD_Displacement.png
  1778   Wed Jul 22 14:44:57 2009 ZachUpdateCamerasGigE Phase Camera

This past week, I have mostly been debugging my software.  I have tried to use the fluorescent lights to test the camera, but I can't tell for sure if my code is finding the correct amplitude and phase or not.  I am currently using Mathematica to double check my calculations in solving for the phase and amplitude.

Also, I have taken dark field images using a lens with a closed shutter.  I have found that the dark band across the top of the images only appears after the camera heats up.  Also, there is an average electronic noise of 14 with a maximum of 40.  However, this electronic noise as well as any consistent ambient noise will be automatically corrected for in the calculations I'm using because I'm taking the differences between the CCD images to calculate relative phases and amplitudes.

I should be able to start setting up optics and performing better tests of my software this week.

  1777   Wed Jul 22 11:18:49 2009 robConfigurationComputerssticky sliders

Quote:

Quote:

Quote:

Quote:

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

 The computers and RFM network are up working again. A boot fest was necessary.Then I restored all the parameters with burtgooey.

The mode cleaner alignment is in a bad state. The autolocker can't get it locked. I don't know what caused it to move so far from the good state that it was till this afternoon.  I went tuning the periscope but the cavity alignment is so bad that it's taking more time than expected. I'll continue working on that tomorrow morning.

 I now suspect that after the reboot the MC mirrors didn't really go back to their original place even if the MC sliders were on the same positions as before.

 Alberto, Rob,

we diagnosed the problem. It was related with sticky sliders. After a reboot of C1:IOO the actual output of the DAC does not correspond anymore to the values read on the sliders. In order to update the actual output it is necessary to do a change of the values of the sliders, i.e. jiggling a bit with them.

 I've updated the slider twiddle script to include the MC alignment biases.  We should run this script whenever we reboot all the hardware, and add any new sticky sliders you find to the end of the script.  It's at

 

/cvs/cds/caltech/scripts/Admin/slider_twiddle

  1776   Wed Jul 22 11:14:11 2009 AlbertoConfigurationComputerscomputers are down

Quote:

Quote:

Quote:

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

 The computers and RFM network are up working again. A boot fest was necessary.Then I restored all the parameters with burtgooey.

The mode cleaner alignment is in a bad state. The autolocker can't get it locked. I don't know what caused it to move so far from the good state that it was till this afternoon.  I went tuning the periscope but the cavity alignment is so bad that it's taking more time than expected. I'll continue working on that tomorrow morning.

 I now suspect that after the reboot the MC mirrors didn't really go back to their original place even if the MC sliders were on the same positions as before.

 Alberto, Rob,

we diagnosed the problem. It was related with sticky sliders. After a reboot of C1:IOO the actual output of the DAC does not correspond anymore to the values read on the sliders. In order to update the actual output it is necessary to do a change of the values of the sliders, i.e. jiggling a bit with them.

  1775   Wed Jul 22 11:08:36 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

I have built a version of the circuit with flying components; the completed circuit is shown in the attached picture. I built the circuit in segments and measured the transfer function after each segment to see whether it matched the LTSpice simulation after each step. The segments are shown in the circuit diagram.

After building the first segment, the measured transfer function looked pretty much the same as the simulated transfer function; it appears shifted in the attached plot, but this is because I didn't do a careful job of tuning at this point, and I'm relatively sure that I could have tuned it to match the simulation. After adding the second segment of the circuit, the measured and simulated transfer functions were similar in shape, but I was unable to increase the frequency of the peaks (through tuning) any more than what is shown in the plot (I could move the peaks so that their frequency was lower, but they are shown as high as they will go). When I added the final segment to complete the circuit, the measured and simulated transfer functions no longer had the same shape; two of the peaks were very close together and I was barely able to differentiate one from the other.

In order to understand what was happening, I tried making modifications to the LTSpice model to recreate the transfer function that was measured. I was able to create a transfer function that closely resembles the measured transfer function in both the circuit as of the 2nd segment and the completed circuit by adding extra inductance and capacitance as shown in red in the circuit diagram. The transfer functions simulated with these parasitic components are shown in red in both plots. While I was able to recreate the response of the circuit, the inductance and capacitance needed to do this were much larger than I would expect to occur naturally within the circuit (2.2uH, 12 pF). I'm not sure what's going on with this.

Attachment 1: BuiltCkt_Picture.png
BuiltCkt_Picture.png
Attachment 2: BuiltCkt2_Final.png
BuiltCkt2_Final.png
Attachment 3: 1stSegment.png
1stSegment.png
Attachment 4: 2ndSegment_ExtraL.png
2ndSegment_ExtraL.png
Attachment 5: Complete_ExtraL.png
Complete_ExtraL.png
  1774   Wed Jul 22 10:49:40 2009 AlbertoUpdatePSLMC Alignment Trend

Trying to track the MC positions back for a few days, it seems that the data hasn't been recorded properly for a while. Something happened yesterday after my boot fest and then the record got restored. Attached here are the readbacks showing the event for MC1.

Is anything wrong with the data record?

Attachment 1: 2009-07-21_MC-align-trend.pdf
2009-07-21_MC-align-trend.pdf
Attachment 2: 2009-07-22_mc-align-trend.pdf
2009-07-22_mc-align-trend.pdf
  1773   Wed Jul 22 09:04:10 2009 AlbertoConfigurationComputerscomputers are down

Quote:

Quote:

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

 The computers and RFM network are up working again. A boot fest was necessary.Then I restored all the parameters with burtgooey.

The mode cleaner alignment is in a bad state. The autolocker can't get it locked. I don't know what caused it to move so far from the good state that it was till this afternoon.  I went tuning the periscope but the cavity alignment is so bad that it's taking more time than expected. I'll continue working on that tomorrow morning.

 I now suspect that after the reboot the MC mirrors didn't really go back to their original place even if the MC sliders were on the same positions as before.

  1772   Wed Jul 22 01:57:19 2009 AlbertoConfigurationComputerscomputers are down

Quote:

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

 The computers and RFM network are up working again. A boot fest was necessary.Then I restored all the parameters with burtgooey.

The mode cleaner alignment is in a bad state. The autolocker can't get it locked. I don't know what caused it to move so far from the good state that it was till this afternoon.  I went tuning the periscope but the cavity alignment is so bad that it's taking more time than expected. I'll continue working on that tomorrow morning.

  1771   Tue Jul 21 18:46:47 2009 steveConfigurationComputerscomputers are down

All suspentions are kicked up. Sus dampings and  oplev servos turned off.

c1iscey and c1lsc are down. c1asc and c1iovme are up-down

  1770   Tue Jul 21 17:52:12 2009 JenneUpdateIOOMC_L flatlined

Quote:

[Clara, Jenne]

While Clara was working on her Wiener filtering and optimizing the locations of the accelerometers, she discovered that MC_L and MC_L_256 are totally flatlined.  I looked at them, and it looks like they've been dead since ~9:30pm-ish on Sunday night.  Bootfest-type activities shall commence shortly.

 Under Alberto's tutalage, I rebooted the whole vme set (iovme, sosvme, susvme1, susvme2), and after that MC_L was all good again.

  1769   Tue Jul 21 17:01:18 2009 peteDAQDAQtemp channel PEM-PETER_FE

I added a temporary channel, to input 9 on the PEM ADCU.    Beware the 30, 31, and 32 inputs.  I tried 32 and it only gave noise.

 

 

  1768   Tue Jul 21 15:32:47 2009 JenneUpdateIOOMC_L flatlined

[Clara, Jenne]

While Clara was working on her Wiener filtering and optimizing the locations of the accelerometers, she discovered that MC_L and MC_L_256 are totally flatlined.  I looked at them, and it looks like they've been dead since ~9:30pm-ish on Sunday night.  Bootfest-type activities shall commence shortly.

  1767   Tue Jul 21 13:55:08 2009 ClaraUpdatePEMGuralp Box Success!

There managed to be just enough 100 kOhm resistors to stuff all the "2" channels (VERT2, N/S2, E/W2) with the fancy low-noise resistors. The first six channels (VERT 1/2, NS 1/2, EW 1/2) are now completely done with the thin-film resistors, taking into account the changes that were made on the circuit diagram. I also replaced the C8 capacitor with the fancy Garrett ones and added capacitors on top of R4 and R13 (after painstakingly making sure that the capacitances are exactly the same for each pair) for the "2" channels. It looks like the capacitors on the "1" channels are the cheaper ones. I will compare the noise measurements later to see if there is any difference - if so, I can replace those as well (although, we're out of the 1 uF capacitors needed for C8).

Speaking of, we are now out of or very low on several types of the Garrett resistors/capacitors: 1 uF, 1kOhm, 100 Ohm, 14.0 Ohm, and 100 kOhm. I left the specifics on Steve's desk so that more can be ordered for the eventual time when the third set of channels needs to be restuffed.

  1766   Tue Jul 21 01:11:30 2009 DmassAoGComputersAlarms going off

I came into the 40m to sign things out briefly then swiftly return them, and the alarms were going off on op540m at 1am.

 

The cat and donkey? were making much noise.

  1765   Mon Jul 20 17:06:29 2009 JenneUpdatePEMGuralp Box Fail

Quote:
That's terrible: R5 & R8 should definitely be 100 Ohm and not 100kOhm. 100k would make it a noise disaster. They should also be metal film (from the expensive box, not from the standard box). This is the same for all channels so might as well stuff them.

The circuit diagram between TP3 and TP4 appears to be designed to make the whitening not work. That's why R6 & R7 should be 100k. And R2 should be metal film too.

Basically, every time we want good low frequency performance we have to use the metal film or metal foil or wirewound resistors. Everything else produces a lot of crackling noise under the influence of DC current.

I'm also attaching the voltage and current noise spectra for the AD620 from the datasheet. These should allow us to compare our measurements to a reasonable baseline.


While we're comparing things to other things, Ben Abbott just emailed me his measurement of the AD620 from back in the day. Clara's going to use this along with the specs to make sure that (a) we're not taking crazy measurements and (b) our AD620s aren't broken and in need of replacement. In this plot, we're looking at the GOLD trace, which has the AD620 set up with a gain of 10, which is how our AD620's are set up in the Guralp breakout box.

Just picking a single point to compare, it looks like at 1Hz, Ben saw ~130dBVrms/rtHz. Converting this to regular units [ 10^(#dB/20)*1Vrms = Vrms ], this is about 3*10^-7 Vrms. That means that Clara's measurements of our AD620 noise is within a factor of 2 of Ben's. Maybe the way we're connecting them up just isn't allowing us to achieve the ~50nV/rtHz that is claimed.
Attachment 1: AD620noise_BenAbbott.pdf
AD620noise_BenAbbott.pdf
  1764   Mon Jul 20 12:35:21 2009 robConfigurationLockingalignment biases funny

I found the alignment biases for the PRM and the SRM in a funny state.  It seemed like they had been "saved" in badly misaligned position, so the restore scripts on the IFO configure screen were not working.  I've manually put them into a better alignment.

  1763   Mon Jul 20 10:35:06 2009 steveUpdateVACUPS batteries replaced

 APC Smart-UPS (uninterruptible power supply) batteries RBC12 replaced at 1Y8 vacuum rack.

Their life span were 22 months.

  1762   Sun Jul 19 22:38:24 2009 robOmnistructureGeneralWeb screenshots aren't being updated

Quote:

Before heading back to the 40m to check on the computer situation, I thought I'd check the web screenshots page that Kakeru worked on, and it looks like none of the screens have been updated since June 1st.  I don't know what the story is on that one, or how to fix it, but it'd be handy if it were fixed.

 Apparently I broke this when I added op540m to the webstatus page.  It's fixed now.

  1761   Sat Jul 18 19:49:48 2009 ranaUpdatePEMGuralp Box Fail
That's terrible: R5 & R8 should definitely be 100 Ohm and not 100kOhm. 100k would make it a noise disaster. They should also be metal film (from the expensive box, not from the standard box). This is the same for all channels so might as well stuff them.

The circuit diagram between TP3 and TP4 appears to be designed to make the whitening not work. That's why R6 & R7 should be 100k. And R2 should be metal film too.

Basically, every time we want good low frequency performance we have to use the metal film or metal foil or wirewound resistors. Everything else produces a lot of crackling noise under the influence of DC current.

I'm also attaching the voltage and current noise spectra for the AD620 from the datasheet. These should allow us to compare our measurements to a reasonable baseline.
Attachment 1: Picture_1.png
Picture_1.png
Attachment 2: Picture_2.png
Picture_2.png
Attachment 3: AD620.pdf
AD620.pdf AD620.pdf AD620.pdf AD620.pdf AD620.pdf AD620.pdf AD620.pdf AD620.pdf
  1760   Fri Jul 17 18:04:54 2009 ClaraUpdatePEMGuralp Box Fail

I've been trying for most of the week to get noise measurements on the output of the Guralp box as well as scross the AD640 chip. The measurements haven't really been making sense, and, being at a loss as to what else I should try, I decided to redo the resistors on the N/S 2 and E/W 2 channels. (I had been comparing the VERT1 and VERT2 channels, as VERT1 has been restuffed and VERT2 has not.) I don't need all three of the second set of channels to do more measurements, so it seemed like a good use of time.

The first thing I noticed was that the VERT2 channel was missing two resistors (R24 and R25). I probably should have noticed this sooner, as they are right by the output points I had been measuring across, but it didn't occur to me that anyone did anything to the VERT2 channel at all. So, probably the measurements on VERT2 are no good.

VERT2_missing_resistors.png

Note the existence of 100 kOhm resistors on the top channel, and none on the bottom channel (VERT2).

 

Then, while I was soldering in some 100 Ohm resistors, I happened to notice that the resistors I was using had a different number (1001) on them than the corresponding ones on the already redone channels (1003). I checked the resistance, and the ones on the already redone channels turned out to be 100 kOhm resistors, rather than 100 Ohms. So, I double checked the circuit diagram to make sure that I had read it correctly, and there were a number of resistors that had been relabeled as 100 Ohms and several relabeled as 100 kOhms. On the board, however, they were ALL 100 kOhms. Clearly, one of them is wrong, and I suspect that it is the circuitboard, but I don't know for sure.

resistors_diagram.png

resistors_board.png

The diagram clearly shows that R6 should be a 100k resistor, while R5 and R8 should be 100 Ohm resistors, but they are all the same (100k) on the board. I suspect this may have something to do with larger-than-expected noise measurements. But, it's possible the diagram is wrong, not the board. In any case, I didn't really know what to do, since I wasn't sure which was right, so I just replaced all the resistors I was sure about and removed the 100k and 100 Ohm resistors without replacing them with anything. Incidentally, the box of 100kOhm resistors seems to be missing, so I wouldn't have been able to finish those anyway.

  1759   Thu Jul 16 14:54:05 2009 robUpdateLockingMC_F channel dead

Quote:

Quote:

It's railed.  This is what halted locking progess on Monday night, as this channel is used for the offloadMCF script, which slowly feeds back a CARM signal to the ETMs to prevent the VCO from saturating.

 

Attached is a 5 day trend, which shows that the channel went dead a few days ago.  All the channels shown are being collected from the same ICS110B (I think), but only some are dead.  It looks like they went dead around the time of the "All computers down" from Sunday.

 Attached are the channels being recorded from the ICS110B in 1Y2 (the IOO rack).  Channels 12, 13, 16, 17, 22, 24, 25 appear to have gone dead after the computer problems on Sunday.

 This has been fixed by one of the two most powerful & useful IFO debugging techniques: rebooting.  I keyed the crate in 1Y2.

  1758   Thu Jul 16 14:41:38 2009 robUpdateLockingMC_F channel dead

Quote:

It's railed.  This is what halted locking progess on Monday night, as this channel is used for the offloadMCF script, which slowly feeds back a CARM signal to the ETMs to prevent the VCO from saturating.

 

Attached is a 5 day trend, which shows that the channel went dead a few days ago.  All the channels shown are being collected from the same ICS110B (I think), but only some are dead.  It looks like they went dead around the time of the "All computers down" from Sunday.

 Attached are the channels being recorded from the ICS110B in 1Y2 (the IOO rack).  Channels 12, 13, 16, 17, 22, 24, 25 appear to have gone dead after the computer problems on Sunday.

Attachment 1: IOO_ICS_0_15.png
IOO_ICS_0_15.png
Attachment 2: IOO_ICS_15_32.png
IOO_ICS_15_32.png
  1757   Thu Jul 16 10:52:58 2009 ClaraUpdatePEMSingle Channel TRS-RNC Cable

I made and tested a female-to-female TRS(audio)-RNC cable. It only has a single channel, so it won't work for stereo speakers or anything, but I should only need one speaker for testing the microphones. The tip of the plug is the signal, the sleeve is ground, and the ring is null.

ELOG V3.1.3-